ADNAN MUHAMMED

Grade 12,

a spring of force constant k is cut into two equal parts,which are then joined parallel to each other.the force constant of the combined will be??? is the answer 'k'.can u verify it?

a spring of force constant k is cut into two equal parts,which are then joined parallel to each other.the force constant of the combined will be??? is the answer 'k'.can u verify it? 

Grade:12

1 Answers

vikas askiitian expert
13 years ago

spring constant of spring of length L is K....

now k is inversly proportional to length ..

  so ,      K2/K1   = L1/L2                          

from here we can compare force constant of two different lengths of  wire of same material..

let force constant of each half is k1 after cutting then

k1/k = L/(L/2)

k1 = 2k

both of these parts are connected parallel so net force constant = 2k+2k=4k

approve my ans if u like it